Solve for a and b
0=a^2 - b^2
1=2ab

Answer by MathPro(14)

Good Afternoon,
Your problem is:
Solve for a and b
0=a^2 - b^2 equation (1)
1=2ab equation (2)
Divide both sides of equation (2) by 2a to isolate b. Note you could also divide it by b instead, it doesn’t matter which variable you choose.
1/(2a) = 2ab/(2a)
1/(2a) = b equation (3)
Now take equation (3) and substitute for b in equation (1) to get:
a^2 - [1/(2a)]^2 = 0 equation (4)
get rid of the fraction by multiplying both sides of equation (4) by (2a)^2 .
a^2 * (2a)^ 2 – 1 = 0(2a)^2 = 0
a^2 * 4a^2 – 1 = 0
4a^2*a^2 – 1 =0
use the product rule: x^a * x^b = x ^ (a+b) to get:
4a^4 – 1 = 0
Add 1 to each side
4a^4 = 1
Divide each side by 4:
a^4 = ¼
use this general rule (x^a)^b = x ^a*b:
raise both sides to the power ¼:
(a^4)^(1/4) = (1/4)^1/4
a^(4*1/4) = (1/4)^ ¼
a = (1/4)^1/4
use the general rule :1 ^ x = 1 to get:
a = (1^1/4) /(4^1/4) = 1/ (4^1/4)
use the general rule: 1/x^a = x^-a to get:
a = 4 ^ -¼ equation (5)
lets rewrite this differently by using 4 = 2*2 = 2^2 to get a =( 2^2)^-1/4 which can be rewritten as
a = 2^2*(-1/4)
a = 2 ^-1/2 equation (5)
from equation (1) we have a^2 = b^2 so b must b the same as a
but let’s double check this by getting b from equation (2)
substitute for a from equation (5) in equation (2) to get b:
1=2*2^-1/2*b
1 = 2 ^ 1 * 2 ^ -1/2 *b
1 = 2^ (1 -1/2) *b by the product rule
1 = 2 ^ ½ *b
Divide both sides by 2^1/2 to get:
b = 1/(2^1/2) = 2^-1/2 equation (6)
Our answer is a = b = 2 ^ -1/2
**Note: we could have also gotten this by inspection. If a=b from equation (1) and ½ =a*b from equation (2) then [a=sqrt(1/2) times b=sqrt(1/2)] could be the only solution that would produce the value ½ in ½ = a*b
Good Luck!
Selma from Rent-A-Whiz, an online tutoring and homework help website. Just visit www.rentawhiz.com to learn more. Currently, we are offering to solve one math problem for free. Just visit our website, select the homework tab, then click on the math problem submission link and quote coupon code 003 when you submit your problem.